# Solving Optimal Power Flow on a Data-Budget: Feature Selection on Smart Meter Data

### v1

@@ III-A OPF Data Compression using PCA
$f_{1}(\mathbf{W})=\frac{1}{2T}\sum_{t=1}^{T}\|\boldsymbol{\theta}_{t}-\mathbf{W}\boldsymbol{\theta}_{t}\|^{2}=\frac{1}{2T}\|\mathbf{\Theta}-\mathbf{W}\mathbf{\Theta}\|_{F}^{2}.$ (12)
@@ III-B OPF Data Selection using DEIM
$\hat{\boldsymbol{\theta}}_{t}=\mathbf{U}_{K}(\mathbf{S}^{\top}\mathbf{U}_{K})^{-1}\tilde{\boldsymbol{\theta}}_{t}=\mathbf{W}_{\text{DEIM}}\boldsymbol{\theta}_{t}$
@@
$f_{1}(\mathbf{W}_{\text{PCA}})\leq f_{1}(\mathbf{W}_{\text{DEIM}})\leq\|(\mathbf{S}^{\top}\mathbf{U}_{K})^{-1}\|^{2}\cdot f_{1}(\mathbf{W}_{\text{PCA}})$
@@ III-C OPF Data Selection using Group Lasso
$\hat{\boldsymbol{\theta}}=\sum_{p=1}^{P}\theta_{p}\mathbf{w}_{p}.$
@@
$\|\mathbf{w}_{n}\|+\|\mathbf{w}_{n+N}\|\quad\longrightarrow\quad\left\|\left[\begin{array}[]{l}\mathbf{w}_{n}\\ \mathbf{w}_{n+N}\end{array}\right]\right\|$
@@
$=\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}}(\mathbf{S}_{\text{GL}}^{\top}\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}})^{-1}.$
@@
$\mathbf{W}_{\text{GL2}}=\mathbf{C}_{\text{GL2}}\mathbf{S}_{\text{GL}}^{\top}=\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}}(\mathbf{S}_{\text{GL}}^{\top}\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}})^{-1}\mathbf{S}_{\text{GL}}^{\top}.$
@@ IV Type-2 OPF Data Distillation
We seek a column-sparse $\mathbf{W}$ so that $\hat{\mathbf{x}}_{t}$ is close to $\mathbf{x}_{t}$ for all $t$. Paralleling (12), we define the OPF fitting error:
@@
Problem (21) is non-convex. Bilevel programs are often handled by replacing every inner problem with its optimality conditions [24]. Unfortunately, such an approach would lead to a large-scale mixed-integer nonlinear program (MINLP) that scales unfavorably with $N$ and $T$. Motivated by advancements in meta-learning, we solve (21) using an APG algorithm. Unfortunately, when it comes to non-convex problems, the standard APG algorithm is not guaranteed to converge to a critical point, that is a point where the cost subdifferential contains zero (zero-gradient point if the cost is differentiable). To resolve this issue, we use a modification of APG proposed in [23, Alg. 2]. Although the modified APG algorithm is slightly more complex, it is guaranteed to converge to a critical point even for non-convex problems. The method is tabulated as Algorithm 2. Its differences to Alg. 1 are discussed next.
@@
$\nabla_{\mathbf{W}}f_{2}(\mathbf{W})=\frac{1}{T}\sum_{t=1}^{T}(\nabla_{\hat{\boldsymbol{\theta}}_{t}}\hat{\mathbf{x}}_{t})^{\top}(\hat{\mathbf{x}}_{t}-\mathbf{x}_{t})\boldsymbol{\theta}_{t}^{\top}.$
@@
$\mathbf{C}_{\text{BGL2}}=\arg\min_{\mathbf{C}}~{}\frac{1}{2T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}(\mathbf{C}\mathbf{S}_{\text{BGL}}^{\top})\|^{2}.$
@@
$\mathbf{W}_{\text{BGL2}}=\mathbf{C}_{\text{BGL2}}\mathbf{S}_{\text{BGL}}^{\top}.$
@@ V Numerical Tests
The proposed OPF data distillation approaches were numerically evaluated using a single-phase rendition of the IEEE 37-bus benchmark feeder and the original, multiphase version of the IEEE 123-bus feeder. Optimal DER reactive power injections $\mathbf{x}_{t}$ and $\hat{\mathbf{x}}_{t}$ were obtained by solving (5) using YALMIP and Sedumi in MATLAB. For the multiphase grid, the objective cost in (5) was modified to $\|q\|^{2}$. All tests were run on a MacBook laptop computer with an M3 Pro chip and 36 GB of RAM. Type-1 methods took less than a few minutes to run. Type-2 methods took from seconds to a few minutes to run for larger $K$ and a few hours for smaller $K$.
V-A Single-phase 37-bus system
@@
Fidelity in OPF data under single-phase system. The first test compares the developed methods in terms of fidelity in reconstructing OPF data. The test involved four Type-1 approaches (PCA, DEIM, GL, and GL2) and one Type-2 approach (BGL). We used the normalized squared error $\tfrac{f_{1}(\mathbf{W})}{\|\mathbf{\Theta}\|_{F}^{2}}$ as a data fidelity metric. This metric was evaluated for an increasing number $K$ of selected OPF features. Data reconstruction errors are expected to decrease for increasing $K$. Parameter $K$ is entered explicitly for PCA and DEIM. For GL-type approaches, we can obtain $K$ non-zero groups in $\mathbf{W}$ by varying parameters $\lambda_{1}$ and $\lambda_{2}$. Figure 2 illustrates the data reconstruction error for varying $K$. PCA serves only as a lower bound as it does not offer data selection. Although DEIM does select data, its performance is unpredictable and exhibits no clear trend. The errors attained by plain GL decrease only marginally as $K$ increases. On the contrary, the GL2 reduces errors more dramatically and follows the trend set by PCA while featuring data selection. BGL behaves similarly to GL and bears no meaningful improvement. Overall, the GL2 approach is the best method to approximate $\mathbf{\Theta}$ when considering sensor selection.
@@
Fidelity in OPF solutions. The second test compares four Type-1 approaches to two Type-2 approaches (BGL and BGL2) in terms of fidelity in finding OPF minimizers $\mathbf{X}$. Figure 3 illustrates the normalized squared error $\tfrac{f_{2}(\mathbf{W})}{\|\mathbf{X}\|_{F}^{2}}$ for increasing $K$. PCA attains the smallest error across all Type-1 designs except for a few cases where GL2 attains smaller errors. DEIM does not show a clear trend and GL shows slight improvement for increasing $K$. Although the errors attained by GL2, BGL and BGL2 decrease when $K$ increases, BGL2 achieves smaller errors. The tests corroborate that BGL2 is the preferred method for approximating OPF solutions $\mathbf{X}$.
@@
Figure 4: Violin plots of voltage magnitudes for the single-phase network experienced under all OPF methods for $K=7$ (top panel) and $K=16$ (bottom panel) OPF features. Violin plots are computed across all buses and time instances. The violin plots obtained using both BGL and two-stage BGL for $K=16$ are almost identical to that of the original OPF method.
@@
Feasibility of OPF solutions. In addition to minimizing $f_{2}(\mathbf{W})$, it is critical that $\hat{\mathbf{x}}_{t}=\mathbf{x}(\hat{\boldsymbol{\theta}}_{t})$ does not violate voltage constraints when applied to the grid under actual loading $\boldsymbol{\theta}_{t}$. We computed the nodal voltages obtained by all OPF data distillation methods according to (2). Regardless of how DER setpoints were obtained, their voltage effect on the grid was computed using the actual loading conditions. Figure 4 depicts the violin plots of all bus voltages across all instances for each method and two values of $K=7$ and $K=16$. Note that even the OPF fed with the original data experiences voltages outside the ±$\pm 3\%$ range because voltage constraints are treated as soft. The tests show that voltage magnitudes are more concentrated under both BGL and BGL2 for both values of $K$. Interestingly, the voltage distributions attained by BGL and BGL2 using only $K=16$ OPF features are almost identical to those achieved by the original OPF fed by all $P=50$ OPF features.
@@
V-B Multiphase 123-bus system
Data generation. We also ran tests on the IEEE 123-bus multiphase distribution feeder [28]. Out of its 262 nodes (combinations of phases and buses excluding the substation), only 95 host non-zero injections. We obtained active power demand and solar generation data at one-minute intervals from the Smart* project [29]. The dataset contained active loads from 444 homes and solar generation from 45 PVs between 5:00 AM and 9:30 PM. We aggregated load demands from multiple homes to simulate 95 active load profiles for the 95 non-zero injection nodes of the tested feeder. Load profiles were scaled to match the kW benchmark load. Reactive load profiles were simulated by drawing power factors uniformly at random within $[0.85,0.99]$ across time and nodes. We also added 95 EV profiles by randomly sampling from [27]. We simulated 10 PVs by aggregating the solar generation from the 45 PVs from [29] and installed them on nodes 2a, 10b, 34a, 46c, 51c, 65a, 76b, 93a, 103b, and 118c.
TABLE I: Fidelity in OPF solution for the IEEE 123-bus system
# selected features $K$ Normalized error in OPF solutions [%]
@@ V-B Multiphase 123-bus system
3 63.62 64.77 65.23 64.72 54.94
26 52.18 64.73 58.10 63.99 31.27
32 46.96 64.72 51.30 63.98 27.59
53 35.03 64.64 42.69 62.80 24.67
Fidelity. The first test for the multiphase network compares GL/GL2 and BGL/BGL2 in terms of reconstructing the OPF minimizers $\mathbf{X}$. Table I shows the normalized squared errors for 4 different values of $K$. Similar to the single-phase network, PCA attains the smallest errors among all Type-1 designs in the multiphase network. Again, BGL2 is the best method for approximating OPF solutions. The reported errors verify the importance of the two-stage approach.
Feasibility. We simulated the approximate OPF solutions along with the true loading conditions to compute nodal voltages per (2). Figure 6 shows the violin plots across all nodes and scenarios for all methods. The tests show that the voltage magnitudes are more concentrated for BGL2 and that by using only $K=53$ OPF features, we can achieve voltage distributions almost identical to those achieved by the original OPF fed by all $P=190$ OPF features.
Figure 6: Violin plots of voltages for the multiphase system for $K=26$ and $53$. BGL2 is almost identical to the original method for $K=53$.

### v2

@@ III-A OPF Data Compression using PCA
$f_{1}(\mathbf{W})$ $=\frac{1}{2T}\sum_{t=1}^{T}\|\boldsymbol{\theta}_{t}-\hat{\boldsymbol{\theta}}_{t}\|^{2}$
$=\frac{1}{2T}\sum_{t=1}^{T}\|\boldsymbol{\theta}_{t}-\mathbf{W}\boldsymbol{\theta}_{t}\|^{2}$
$=\frac{1}{2T}\|\mathbf{\Theta}-\mathbf{W}\mathbf{\Theta}\|_{F}^{2}.$ (12)
@@ III-B OPF Data Selection using DEIM
$\hat{\boldsymbol{\theta}}_{t}=\mathbf{U}_{K}(\mathbf{S}^{\top}\mathbf{U}_{K})^{-1}\tilde{\boldsymbol{\theta}}_{t}=\mathbf{W}_{\text{DEIM}}\boldsymbol{\theta}_{t}$ (14)
@@
$f_{1}(\mathbf{W}_{\text{PCA}})\leq f_{1}(\mathbf{W}_{\text{DEIM}})\leq\|(\mathbf{S}^{\top}\mathbf{U}_{K})^{-1}\|^{2}f_{1}(\mathbf{W}_{\text{PCA}})$ (15)
@@ III-C OPF Data Selection using Group Lasso
$\hat{\boldsymbol{\theta}}=\sum_{p=1}^{P}\theta_{p}\mathbf{w}_{p}.$ (17)
@@
$\|\mathbf{w}_{n}\|+\|\mathbf{w}_{n+N}\|\quad\longrightarrow\quad\left\|\left[\begin{array}[]{l}\mathbf{w}_{n}\\ \mathbf{w}_{n+N}\end{array}\right]\right\|$ (18)
@@
$=\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}}(\mathbf{S}_{\text{GL}}^{\top}\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}})^{-1}.$ (23)
@@
$\mathbf{W}_{\text{GL2}}=\mathbf{C}_{\text{GL2}}\mathbf{S}_{\text{GL}}^{\top}=\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}}(\mathbf{S}_{\text{GL}}^{\top}\mathbf{C}_{\theta}\mathbf{S}_{\text{GL}})^{-1}\mathbf{S}_{\text{GL}}^{\top}.$ (24)
@@ IV Type-2 OPF Data Distillation
We seek a column-sparse $\mathbf{W}$ so that $\hat{\mathbf{x}}_{t}$ is close to $\mathbf{x}_{t}$ for all $t$. Paralleling (III-A), we define the OPF fitting error:
@@
Problem (27) is non-convex. Bilevel programs are often handled by replacing every inner problem with its optimality conditions [31]. Unfortunately, such an approach would lead to a large-scale mixed-integer nonlinear program (MINLP) that scales unfavorably with $N$ and $T$. Motivated by advancements in meta-learning, we solve (27) using an APG algorithm. Unfortunately, when it comes to non-convex problems, the standard APG algorithm is not guaranteed to converge to a critical point, that is a point where the cost subdifferential contains zero (zero-gradient point if the cost is differentiable). To resolve this issue, we use a modification of APG proposed in [30, Alg. 2]. Although the modified APG algorithm is slightly more complex, it is guaranteed to converge to a critical point even for non-convex problems. The method is tabulated as Algorithm 2. Its differences to Alg. 1 are discussed next.
@@
$\nabla_{\mathbf{W}}f_{2}(\mathbf{W})=\frac{1}{T}\sum_{t=1}^{T}(\nabla_{\hat{\boldsymbol{\theta}}_{t}}\hat{\mathbf{x}}_{t})^{\top}(\hat{\mathbf{x}}_{t}-\mathbf{x}_{t})\boldsymbol{\theta}_{t}^{\top}.$ (28)
@@
$\mathbf{C}_{\text{BGL2}}=\arg\min_{\mathbf{C}}~{}\frac{1}{2T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}(\mathbf{C}\mathbf{S}_{\text{BGL}}^{\top})\|^{2}.$ (29)
@@
$\mathbf{W}_{\text{BGL2}}=\mathbf{C}_{\text{BGL2}}\mathbf{S}_{\text{BGL}}^{\top}.$ (30)
@@ V Numerical Tests
The proposed OPF data distillation approaches were numerically evaluated using a single-phase rendition of the IEEE 37-bus benchmark feeder and a single-phase 1,136-bus feeder obtained from the IEEE 8,500-bus feeder [26]. Optimal DER reactive power injections $\mathbf{x}_{t}$ and $\hat{\mathbf{x}}_{t}$ were obtained by solving (5) using YALMIP and SeDuMi in MATLAB [33, 34]. All tests were run on a MacBook laptop computer with an M3 Pro chip and 36 GB of RAM. Type-1 methods took less than a few minutes to run. Type-2 methods took from seconds to a few minutes to run for smaller $K$, and a few hours for larger $K$.
V-A Numerical Tests on the 37-bus Feeder
@@
Fidelity in OPF data under single-phase system. The first test compares the proposed methods in terms of fidelity in reconstructing OPF data. The test involved four Type-1 approaches (PCA, DEIM, GL, and GL2) and two Type-2 approaches (BGL). We used the normalized squared error $\tfrac{f_{1}(\mathbf{W})}{\|\mathbf{\Theta}\|_{F}^{2}}$ as a data fidelity metric. This metric was evaluated for an increasing number $K$ of selected OPF features. Data reconstruction errors are expected to decrease for increasing $K$. Parameter $K$ is entered explicitly for PCA and DEIM. For GL-type approaches, we can obtain $K$ non-zero groups in $\mathbf{W}$ by varying parameters $\lambda_{1}$ and $\lambda_{2}$. Figure 2 illustrates the data reconstruction error for varying $K$. PCA serves only as a lower bound, as it does not offer data selection. Although DEIM does select data, its performance is unpredictable and exhibits no clear trend. The errors attained by plain GL decrease only marginally as $K$ increases. On the contrary, the GL2 reduces errors more dramatically and follows the trend set by PCA while featuring data selection. BGL behaves similarly to GL and bears no meaningful improvement. Not shown in Figure 2 is the performance curve of BGL2. BGL2 had the worst data reconstruction errors among all methods for most values of $K$, and the errors did not follow any specific trend. Overall, the GL2 approach is the most effective method for approximating $\mathbf{\Theta}$ when considering sensor selection.
@@
Fidelity in OPF solutions. The second test compares four Type-1 approaches to two Type-2 approaches (BGL and BGL2) in terms of fidelity in finding OPF minimizers $\mathbf{X}$. Figure 3 illustrates the normalized squared error $\tfrac{f_{2}(\mathbf{W})}{\|\mathbf{X}\|_{F}^{2}}$ for increasing $K$. PCA achieves the smallest error across all Type-1 designs, except in a few cases where GL2 yields smaller errors. DEIM does not exhibit a clear trend, and GL shows a slight improvement with increasing $K$. Although the errors attained by GL2, BGL, and BGL2 decrease when $K$ increases, BGL2 achieves smaller errors. The tests corroborate that BGL2 is the preferred method for approximating OPF solutions. From Figure 3, it can be seen that when $K\geq 25$, the minimizer errors of GL2 are very close to those obtained by BGL and BGL2. At $K=31$, GL2 achieves a minimizer error of about $6\%$, whereas BGL and BGL2 achieve minimizer errors of about $3\%$ and $1\%$, respectively. Therefore, at larger values of $K$, GL2 may be preferable, as it can achieve small minimization errors while forgoing the increased computational burden of BGL or BGL2. This behavior is expected as discussed in Remark 1.
@@
Figure 4: Violin plots of voltage magnitudes computed using the linearized grid model of (2) for the IEEE 37-bus single-phase network for $K=7$ (top panel) and $K=16$ (bottom panel) OPF features. Voltages are computed across all buses and scenarios using the actual OPF data. The two black solid horizontal lines mark the ±$\pm 3\%$ deviations in voltages. The violin plots obtained using both BGL and two-stage BGL for $K=16$ are almost identical to those of the original OPF method.
Figure 5: Violin plots of voltage magnitudes computed using the exact AC grid model for the IEEE 37-bus single-phase network for $K=7$ (top panel) and $K=16$ (bottom panel) OPF features. Voltages are computed across all buses and scenarios using the actual OPF data. The two black solid horizontal lines mark the ±$\pm 3\%$ deviations in voltages. The violin plots obtained using both BGL and two-stage BGL for $K=16$ are almost identical to those of the original OPF method.
@@
Feasibility of OPF solutions. In addition to minimizing $f_{2}(\mathbf{W})$, it is critical that $\hat{\mathbf{x}}_{t}=\mathbf{x}(\hat{\boldsymbol{\theta}}_{t})$ does not violate voltage constraints when applied to the grid under actual loading $\boldsymbol{\theta}_{t}$. We computed the nodal voltages obtained by all OPF data distillation methods according to the linearized grid model of (2). Regardless of how DER setpoints were obtained, their voltage effect on the grid was computed using the actual loading conditions. Figure 4 depicts the violin plots of all bus voltages across all instances for each method and two values of $K=7$ and $K=16$. Note that even the OPF fed with the original data experiences voltages outside the ±$\pm 3\%$ range because voltage constraints are treated as soft constraints. The tests show that voltage magnitudes are more concentrated under both BGL and BGL2 for both values of $K$, with a few extreme outliers. Interestingly, the voltage distributions attained by BGL2 using only $K=16$ OPF features are almost identical to those achieved by the original OPF fed by all $P=50$ features.
Figure 5 depicts violin plots of voltages based on the exact AC grid model. The obtained voltage distributions are similar to those of Fig. 4, and again, BGL2 with $K=16$ provides voltage distributions almost identical to those achieved by the original OPF fed by all $P=50$ OPF features. If the focus is on voltage feasibility rather than fidelity in OPF decisions, BGL/BGL2 may exhibit no significant advantage for larger values of $K$.
@@
Sample size effect. We next evaluated how using different sample sizes affected our proposed methods. Fifty (50) scenarios were randomly sampled from the $T=800$ scenarios that were used in previous tests. The same $\lambda_{2}$ values resulted in different values of $K$. For the same number of selected features $K=5$, different features were selected for the two scenarios. The minimizer errors for all Type-1 and Type-2 methods are lower when the number of OPF scenarios is smaller. Data reconstruction errors are also lower when the OPF data has fewer scenarios. The tests show that the number of OPF scenarios affects both data and minimizer reconstruction errors and also decides which OPF features are selected.
V-B Numerical Tests on the 1,136-bus Feeder
Data generation. To test the scalability of the proposed methods, we also ran tests on a single-phase 1,136-bus feeder obtained from the IEEE 8,500-bus test feeder by keeping buses sited below bus 338 [26]. Out of its 1,136 nodes, only 519 host non-zero injections. We obtained active power demand and solar generation data at one-minute intervals from the Smart* project [37]. The dataset contained active loads from 444 homes and solar generation from 45 PVs between 5:00 AM and 9:30 PM. We aggregated load demands from multiple homes to simulate 519 active load profiles for the 519 non-zero injection nodes of the tested feeder. Load profiles were scaled to match the kW benchmark load. Reactive load profiles were simulated by drawing power factors uniformly at random within $[0.85,0.99]$ across time and nodes. We simulated 20 PVs by aggregating the solar generation from the 45 PVs from [37] and randomly installed them on 20 of the non-zero injection nodes (514, 535, 589, 600, 601, 616, 656, 664, 707, 730, 781, 839, 840, 897, 914, 959, 981, 1036, 1058 and 1069).
TABLE I: Fidelity in OPF solutions for the IEEE 1,136-bus system
$K$ out of $P=1,038$ Normalized error in OPF solutions [%]
@@ V-B Numerical Tests on the 1,136-bus Feeder
2 99.29 99.99 99.65 88.55 40.22
5 98.29 99.99 99.11 53.54 36.30
21 93.28 99.96 96.33 39.95 28.35
36 88.86 99.92 93.80 20.69 6.02
116 68.98 99.77 80.91 18.11 4.91
Fidelity. The first test for the 1,136-bus network compares GL/GL2 and BGL/BGL2 in terms of reconstructing the OPF minimizers $\mathbf{X}$. Table I shows the normalized squared errors for 5 values of $K$. Similar to the smaller 37-bus network, PCA attains the smallest errors among all Type-1 designs in the 1,136-bus network. Again, BGL2 is the best method for approximating OPF solutions. The reported errors verify the importance of the two-stage approach.
Feasibility. We simulated the approximate OPF solutions along with the true loading conditions to compute nodal voltages per (2). Figure 7 shows the violin plots across all nodes and scenarios for all methods. The tests show that the voltage magnitudes are more concentrated for BGL2 and that by using only $K=116$ OPF features, we can achieve voltage distributions almost identical to those achieved by the original OPF fed by all $P=1,038$ OPF features. Figure 8 depicts violin plots of voltages based on the exact AC grid model. Conclusions similar to those of Fig. 7 can be drawn.
Figure 7: Violin plots of voltage magnitudes for the single-phase 1,136-bus network experienced under all OPF methods for $K=21$ (top) and $K=116$ (bottom) out of 1,038 OPF features. Voltages are computed across all buses and scenarios using the linearized grid model of (2) and actual OPF data. The two black solid horizontal lines mark the ±$\pm 3\%$ deviations in voltages. The violin plots obtained using BGL2 with $K=116$ are the closest to those obtained from the original OPF method.
Figure 8: Violin plots of voltage magnitudes computed using the exact AC grid model for the single-phase 1,136-bus network experienced under all OPF methods for $K=21$ (top) and $K=116$ (bottom) out of 1,038 OPF features. Voltages are computed across all buses and scenarios using the linearized grid model of (2) and actual OPF data. The violin plots obtained using BGL2 with $K=116$ are the closest to those obtained from the original OPF method.